# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### 18 Dec 2023 v2

@@ A.1.1 Add Gaussian Noise to Original Images
$\frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}=\frac{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{X_{i}}}\mathrm{cov}^{2}(X_{i},Y_{i}))}{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},X_{i})\frac{1}{\sigma^{2}_{X_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))})}$ (56)
therefore, the entropy change is expressed as:
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&\frac{1}{2}\log\frac{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{X_{i}}}\mathrm{cov}^{2}(X_{i},Y_{i}))}{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},X_{i})\frac{1}{\sigma^{2}_{X_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))})}\\ =&\frac{1}{2}\log\frac{1}{1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}}\end{split}$ (57)
It is difficult to tell that Equation 56 is greater or smaller than $1$ directly. But one thing for sure is that when there is no Gaussian noise, Equation 56 equals 1. However, we can use another way to compare the numerator and denominator of Equation 56. Instead, we compare the numerator and denominator using subtraction. Let:
$\begin{split}f(\sigma_{\epsilon}^{2})=&1-({1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}})\\ =&-\lambda{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}}\\ =&-\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{X_{i}}}}{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}}\end{split}$ (58)
Obviously, the variance $\sigma_{\epsilon}^{2}$ of the Gaussian noise control the result of $f(\sigma_{\epsilon})$, while the mean $\mu_{\epsilon}$ has no influence. When $\sigma_{\epsilon}$ approaching 0, we have:
$\lim_{\sigma_{\epsilon}^{2}\to 0}f(\sigma_{\epsilon}^{2})=0$ (59)
To determine if Gaussian noise can be positive noise, we need to determine whether the entropy change is large or smaller than 0:
$\left\{\begin{matrix}\frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}>1&\mathrm{if}\ f(\sigma_{\epsilon}^{2})>0\\ \frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}\leq 1&\mathrm{if}\ f(\sigma_{\epsilon}^{2})\leq 0\end{matrix}\right.$ (60)
Combine the Equations 43 and 60, we can get the conclusion:
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})>0&\mathrm{if}\ f(\sigma_{\epsilon}^{2})>0\\ \bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})\leq 0&\mathrm{if}\ f(\sigma_{\epsilon}^{2})\leq 0\end{matrix}\right.$ (61)
From the above equations, the sign of the entropy change is determined by the statistical properties of the data samples and labels. Since ${\epsilon}^{2}\geq 0$ and $\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{X_{i}}}\geq 0$, we have a deep dive into the residual part, i.e.,
${\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}}=\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{4}_{X_{i}}\sigma^{2}_{Y_{i}}(1-\rho_{X_{i}Y_{i}}^{2})}$ (62)
where $\rho_{X_{i}Y_{i}}$ is the correlation coefficient, and $\rho^{2}_{X_{i}Y_{i}}\in[0,1]$. As a result, the sign of the entropy change in the Gaussian noise case is negative. We can conclude that Gaussian noise added to the images is harmful to the task.
A.1.2 Inject Gaussian Noise in Latent Space
In this case, the task entropy is formulated as:
$\begin{matrix}H(\mathcal{T};\boldsymbol{X}+\boldsymbol{\epsilon})=H(\boldsymbol{Y};\boldsymbol{X}+\boldsymbol{\epsilon})-H(\boldsymbol{X})\end{matrix}.$ (63)
Thus, the entropy change of injecting Gaussian noise can be formulated as:
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{X})-(H(\boldsymbol{Y};\boldsymbol{X}+\boldsymbol{\epsilon})-H(\boldsymbol{X}))\\ =&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{Y};\boldsymbol{X}+\boldsymbol{\epsilon})\end{split}$ (64)
Borrow the equations from the case of Gaussian noise added the original image, we have:
@@ A.1.2 Inject Gaussian Noise in Latent Space
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{Y};\boldsymbol{X}+\boldsymbol{\epsilon})\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{X}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}|}{|\boldsymbol{\Sigma_{X+\epsilon}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{XY}}|}\\ =&\frac{1}{2}\log\frac{1}{(1+\sigma^{2}_{\epsilon}{\textstyle\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{X_{i}}}})(1+\lambda{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}})}\end{split}$ (65)
Obviously, injecting Gaussian noise into the latent space is harmful to the task.
A.2 Influence of Linear Transform Noise on Task Entropy
In our work, the linear transform noise refers to an image or the latent representation of an image that is perturbed by the combination of other images or latent representations of other images.
A.2.1 Add Linear Transform Noise to Original Images
The task entropy with linear transform noise can be formulated as:
where $I$ is an identity matrix, and $Q$ is derived from $I$ using elementary row operations. The conditional distribution of $\boldsymbol{Y}$ given $\boldsymbol{X}+Q\boldsymbol{X}$ is also multivariate subjected to the normal distribution, which can be formulated as:
The linear transform on $\boldsymbol{X}$ does not change the distribution of the $\boldsymbol{X}$. It is not difficult to obtain:
The entropy change of 0 indicates that the implementation of linear transformation to the original images could not reduce the complexity of the task.
A.2.2 Inject Linear Transform Noise in Latent Space
The entropy change of injecting linear transform noise can be formulated as:
@@ A.2.2 Inject Linear Transform Noise in Latent Space
$\begin{split}\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})=&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{X})-(H(\boldsymbol{Y};\boldsymbol{X}+Q\boldsymbol{X})-H(\boldsymbol{X}))\\ =&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{Y};\boldsymbol{X}+Q\boldsymbol{X})\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma}_{\boldsymbol{X}}||\boldsymbol{\Sigma_{Y}}-\boldsymbol{\Sigma_{YX}}{\boldsymbol{\Sigma_{X}^{-1}}}\boldsymbol{\Sigma_{XY}}|}{|\boldsymbol{\Sigma}_{(I+Q)\boldsymbol{X}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YX}}\boldsymbol{\Sigma^{-1}_{X}}\boldsymbol{\Sigma_{XY}}|}\\ =&\frac{1}{2}\log\frac{1}{|I+Q|^{2}}\\ =&-\log|I+Q|\end{split}$ (71)
@@
$\begin{split}&\max_{Q}\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})\\ &s.t.\ rank(I+Q)=k\\ &\ \ \quad Q\sim I\\ &\ \quad\left[I+Q\right]_{ii}\geq\left[I+Q\right]_{ij},i\neq j\\ &\ \quad\left\|\left[I+Q\right]_{i}\right\|_{1}=1\end{split}$ (72)
where $\sim$ means the row equivalence. The key to determining whether the linear transform is positive noise or not lies in the matrix of $Q$. The most important step is to ensure that $I+Q$ is reversible, which is $|(I+Q)|\neq 0$. For this, we need to investigate what leads $I+Q$ to be rank-deficient. The third constraint is to make the trained classifier get enough information about a specific image and correctly predict the corresponding label. For example, for an image $X_{1}$ perturbed by another image $X_{2}$, the classifier obtained dominant information from $X_{1}$ so that it can predict the label $Y_{1}$. However, if the perturbed image $X_{2}$ is dominant, the classifier can hardly predict the correct label $Y_{1}$. The fourth constraint is the normalization of latent representations.
@@
Full Rank Cases Except for the rank deficiency cases, $I+Q$ has full rank and is reversible. Since $Q$ is a row equivalent to the identity matrix, we need to introduce the three types of elementary row operations as follows [51].
@@
    Row swap here allows exchanging any number of rows. This is slightly different from the original one that only allows any two rows exchange since following the original row swap will lead to a rank deficiency. When the $Q$ is derived from $I$ with Row Swap, it will break the third constraint. Therefore, Row Swap merely is considered harmful and would degrade the deep model.
@@
    Scalar Multiplication Multiply any row by a constant $\beta$. This breaks the fourth constraint, thus degrading the deep models.
@@
    where $\beta$ can be at a random position beside the diagonal. As we can see from the simple example, Row Sum breaks the fourth constraint and make entropy change smaller than 0.
@@
However, if we combine the elementary row operations, it is possible to make the entropy change greater than 0 as well as satisfy the constraints. For example, we combine the Row Swap and Scalar Multiplication to generate the $Q$:
@@
In this case, $\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})>0$ when $Q=-0.5I$. The constraints are satisfied. This is just a simple case that adding linear transform noise that benefits deep models. Actually, there exists a design space of $Q$ that within the design space, deep models can reduce task entropy by injecting linear transform noise. To this end, we demonstrate that linear transform can be positive noise.
A.3 Influence of Salt-and-pepper Noise on Task Entropy
A.3.1 Add Salt-and-pepper Noise to Original Images
The task entropy with salt-and-pepper noise is rewritten as:
where
Therefore, the entropy change with salt-and-pepper noise is:
$\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=H(\mathcal{T};\boldsymbol{X})-H(\mathcal{T};\boldsymbol{X}\boldsymbol{\epsilon})=0$ (82)
A.3.2 Add Salt-and-pepper Noise in Latent Space
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{X})-(H(\boldsymbol{Y};\boldsymbol{X}\boldsymbol{\epsilon})-H(\boldsymbol{X}))\\ =&H(\boldsymbol{Y};\boldsymbol{X})-H(\boldsymbol{Y};\boldsymbol{X}\boldsymbol{\epsilon})\\ =&-\sum_{\boldsymbol{X}\in\mathcal{X}}\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{X},\boldsymbol{Y})\log p(\boldsymbol{X},\boldsymbol{Y})-\sum_{\boldsymbol{X}\in\mathcal{X}}\sum_{\boldsymbol{Y}\in\mathcal{Y}}\sum_{\boldsymbol{\epsilon}\in\mathcal{E}}p(\boldsymbol{X\epsilon},\boldsymbol{Y})\log p(\boldsymbol{X\epsilon},\boldsymbol{Y})\\ =&\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{X},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{X\epsilon},\boldsymbol{Y})}\right]\\ =&\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{X},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{X},\boldsymbol{Y})}\right]-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{\epsilon})}\right]\\ =&-\mathbb{E}\left[\log\frac{1}{p(\boldsymbol{\epsilon})}\right]\\ =&-H(\epsilon)\end{split}$ (83)
The entropy change is smaller than 0, therefore, the salt-and-pepper is a pure detrimental noise to the latent representations.
Appendix B Optimal Quality Matrix of Linear Transform Noise
The optimal quality matrix should maximize the entropy change, therefore theoretically define the minimized task complexity. The optimization problem as formulated in Equation 18 is:

### 02 Feb 2024 v3

@@ B.1 Inject Gaussian Noise in Latent Space
$\frac{|\boldsymbol{\Sigma_{Y|Z}}|}{|\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}=\frac{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\frac{1}{\sigma^{2}_{Z_{i}}}\mathrm{cov}^{2}(Z_{i},Y_{i}))}{\prod_{i=1}^{k}(\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Y_{i},Z_{i})\frac{1}{\sigma^{2}_{Z_{i}}})\cdot(1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))})}$ (60)
Combine Equations 60 and 50 together, the entropy change is expressed as:
@@
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&\frac{1}{2}\log\frac{1}{(1+\sigma^{2}_{\epsilon}{\textstyle\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}})(1+\lambda{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))}})}\end{split}$ (61)
It is difficult to tell that Equation 61 is greater or smaller than 00 directly. But one thing for sure is that when there is no Gaussian noise, Equation 61 equals 0. However, we can use another way to compare the numerator and denominator in Equation 61. Instead, we use the symbol $M$ to compare the numerator and denominator using subtraction. Let:
$\begin{split}M=&1-(1+\sigma^{2}_{\epsilon}{\textstyle\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}})(1+\lambda{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))}})\end{split}$ (62)
Obviously, the variance $\sigma_{\epsilon}^{2}$ of the Gaussian noise control the result of $M$, while the mean $\mu_{\epsilon}$ has no influence. When $\sigma_{\epsilon}$ approaching 0, we have:
$\lim_{\sigma_{\epsilon}^{2}\to 0}M=0$ (63)
To determine if Gaussian noise can be positive noise, we need to determine whether the entropy change is large or smaller than 0.
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})>0&\mathrm{if}\ M>0\\ \bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})\leq 0&\mathrm{if}\ M\leq 0\end{matrix}\right.$ (64)
From the above equations, the sign of the entropy change is determined by the statistical properties of the embeddings and labels. Since ${\epsilon}^{2}\geq 0$, $\lambda\geq 0$ and $\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}\geq 0$, we need to have a deep dive into the residual part, i.e.,
${\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{Z_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))}}=\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{4}_{Z_{i}}\sigma^{2}_{Y_{i}}(1-\rho_{Z_{i}Y_{i}}^{2})}$ (65)
where $\rho_{Z_{i}Y_{i}}$ is the correlation coefficient, and $\rho^{2}_{Z_{i}Y_{i}}\in[0,1]$. Eq. 65 is greater than 0, As a result, the sign of the entropy change in the Gaussian noise case is negative. We can conclude that Gaussian noise added to the latent space is harmful to the task.
B.2 Add Gaussian Noise to Raw Images
Assuming that the pixels of the raw images follow a Gaussian distribution. The variation of task complexity by adding Gaussian noise to raw images can be formulated as:
Borrow the equations from the case of Gaussian noise added to the latent space, we have:
@@ B.2 Add Gaussian Noise to Raw Images
$\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=\frac{1}{2}\log\frac{1}{1+\lambda\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(X_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{X_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(X_{i},Y_{i}))}}$ (67)
Clearly, the introduction of Gaussian noise to each pixel in the original images has a detrimental impact on the task. Note that some studies have empirically shown that adding Gaussian noise to partial pixels of input images may be beneficial to the learning task [30] [69].
Appendix C Impact of Linear Transform Noise on Task Entropy
In our work, concerning the image level perspective, "linear transform noise" denotes an image that is perturbed by another image or a combination of other images. From the viewpoint of embeddings, "linear transform noise" refers to an embedding perturbed by another embedding or the combination of other embeddings.
C.1 Inject Linear Transform Noise in Latent Space
The entropy change of injecting linear transform noise into embeddings can be formulated as:
$\begin{split}\bigtriangleup S(\mathcal{T},Q\boldsymbol{Z})=&H(\mathcal{T};\boldsymbol{Z})-H(\mathcal{T};\boldsymbol{Z}+Q\boldsymbol{Z})\\ =&H(\boldsymbol{Y};\boldsymbol{Z})-H(\boldsymbol{Z})-(H(\boldsymbol{Y};\boldsymbol{Z}+Q\boldsymbol{Z})-H(\boldsymbol{Z}))\\ =&H(\boldsymbol{Y};\boldsymbol{Z})-H(\boldsymbol{Y};\boldsymbol{Z}+Q\boldsymbol{Z})\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma}_{\boldsymbol{Z}}||\boldsymbol{\Sigma_{Y}}-\boldsymbol{\Sigma_{YZ}}{\boldsymbol{\Sigma_{Z}^{-1}}}\boldsymbol{\Sigma_{ZY}}|}{|\boldsymbol{\Sigma}_{(I+Q)\boldsymbol{Z}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}}\boldsymbol{\Sigma^{-1}_{Z}}\boldsymbol{\Sigma_{ZY}}|}\\ =&\frac{1}{2}\log\frac{1}{|I+Q|^{2}}\\ =&-\log|I+Q|\end{split}$ (68)
@@ C.1 Inject Linear Transform Noise in Latent Space
$\begin{split}&\max_{Q}\bigtriangleup S(\mathcal{T},Q\boldsymbol{Z})\\ &s.t.\ rank(I+Q)=k\\ &\ \ \quad Q\sim I\\ &\ \quad\left[I+Q\right]_{ii}\geq\left[I+Q\right]_{ij},i\neq j\\ &\ \quad\left\|\left[I+Q\right]_{i}\right\|_{1}=1\end{split}$ (69)
where $\sim$ means the row equivalence. The key to determining whether the linear transform is positive noise or not lies in the matrix of $Q$. The most important step is to ensure that $I+Q$ is invertible, which is $|(I+Q)|\neq 0$. For this, we need to investigate what leads $I+Q$ to be rank-deficient. The third constraint is to make the trained classifier get enough information about a specific embedding of an image and correctly predict the corresponding label. For instance, when an embedding $Z_{1}$ is perturbed by another embedding $Z_{2}$, the classifier predominantly relies on the information from $Z_{1}$ to predict the label $Y_{1}$. Conversely, if the perturbed embedding $Z_{2}$ takes precedence, the classifier struggles to accurately predict the label $Y_{1}$ and is more likely to predict it as label $Y_{2}$. The fourth constraint is the normalization of latent representations.
@@
Full Rank Cases Except for the rank deficiency cases, $I+Q$ has full rank and is invertible. Since $Q$ is a row equivalent to the identity matrix, we need to introduce the three types of elementary row operations as follows [52].
@@
    Row swap here allows exchanging any number of rows. This is slightly different from the original one that only allows any two-row exchange since following the original row swap will lead to a rank deficiency. When the $Q$ is derived from $I$ with Row Swap, it will break the third constraint. Therefore, Row Swap merely is considered harmful and would degrade the performance of deep models.
@@
    Scalar Multiplication Multiply any row by a constant $\beta$. This breaks the fourth constraint, thus degrading the performance of deep models.
@@
    where $\beta$ can be at a random position beside the diagonal. As we can see from the simple example, Row Sum breaks the fourth constraint and makes entropy change smaller than 0.
@@
However, if we combine the elementary row operations, it is possible to make the entropy change greater than 0 as well as satisfy the constraints. For example, we combine the Row Sum and Scalar Multiplication to generate the $Q$:
@@
In this case, $\bigtriangleup S(\mathcal{T},Q\boldsymbol{X})>0$ when $Q=-0.5I$. The constraints are satisfied. This is just a simple case of adding linear transform noise that benefits deep models. Actually, there exists a design space of $Q$ that within the design space, deep models can reduce task entropy by injecting linear transform noise into the embeddings. To this end, we demonstrate that linear transform can be positive noise.
From the discussion in this section, we can draw conclusions that Linear Transform Noise can be positive under certain conditions, while Gaussian Noise and Salt-and-pepper Noise are harmful noise. From the above analysis, the conditions that satisfy positive noise form a design space. Exploring the design space of positive noise is an important topic for future work.
C.1.1 Optimal Quality Matrix of Linear Transform Noise
The optimal quality matrix should maximize the entropy change and therefore theoretically define the minimized task complexity. The optimization problem as formulated in Equation 19 is: